OpenStudy (anonymous):

Need help with factoring. 5x^.5+x+4

OpenStudy (akashdeepdeb):

\[5\sqrt{x} + x + 4\] \[Take ~~\sqrt{x} = y~~~\therefore ~~x = y^2\] We do this so that we can get rid of the radical and we can factor it easily and then substitute \(\sqrt{x}\) back in the factored polynomial. Our polynomial now becomes: \(5y + y^2 + 4 = y^2 + 5y + 4\) Can you factorize this now? :)

OpenStudy (anonymous):

(y+4)(y+1)? But I still dont see how this gives me the factor of the inital equation.

OpenStudy (anonymous):

Unless I replace the y with square root of x?

OpenStudy (anonymous):

I don't know.

OpenStudy (jhannybean):

we can rewrite this as \(\ x+5\sqrt{x} +4\) so all we have to think about is what multiplies to give us 4 and adds to give us 5?

OpenStudy (anonymous):

4 and 1?

OpenStudy (jhannybean):

Yep, and you know that \(\ x = \sqrt{x} \cdot \sqrt{x} \)

OpenStudy (jhannybean):

Therefore you can factor it out as : \(\ (\sqrt{x} +1)(\sqrt{x}+4)\)

OpenStudy (anonymous):

Oh, I get it now. And I see where Aka was going with this earlier.

OpenStudy (freckles):

You are exactly right to replace that y up there with sqrt(x)

OpenStudy (anonymous):

Thanks for all your help :)

OpenStudy (jhannybean):

Yeah, me and him had 2 different approaches :P